# Finite Difference Solution Ansatz approach in Least-Squares Monte Carlo

### v8

@@ 2.3.2 FD Solution: Control Variate
  • As mentioned before, $\hat{V}_{k+1}^{1D}$ can be seen as option value under a 1D simulation process. However, instead of employing the right-hand side in Eqn.2.32 directly for RSS reduction, Prop.1 allows one to achieve exactly the same effect by just adding the FD ansatz into the projection basis functions. This helps mitigate computational cost as it saves additional tedious Monte Carlo simulation on $\hat{V}^{1D}$, as well as the variance and covariance calculation in Eqn.2.23.
  • From the proof of Prop.1, up to first order of RSS, we don’t see benefit of introducing cross term between $\vec{\phi}^{P^{\text{FD}}}(x)$ and $\vec{\phi}_{R-1}^{P^{\text{LSM}}}(x)$. This simplifies the construction of new basis and avoid overfitting due to further introduction of higher order terms.
  • The most effective FD auxiliary variable $\hat{V}^{1D}$ is obtained by achieving the highest possible correlation with $V$ based on Eqn.2.33. This provides an important theoretical foundation on the choice of FD ansatz in FD-LSM. For instance, under a multi-asset single-factor setting (e.g. Black Scholes or Local Volatility), we can choose the ansatz from single-factor 1D PDE after inter-asset dimension reduction; for single-asset multi-factor settings (e.g. Heston stochastic volatility), one can construct the equivalent 1D process (Local volatility in general) by matching same marginal distributions as mentioned before, see Gyöngy, (1986); finally, for the most general case of multi-asset multi-factor setting, one could apply the dimension reduction on intra-asset level (reducing the factors other than spot, e.g. stochastic volatility on the same asset) to start with, and then on inter-asset level, eventually arriving at 1D problem which is solvable under Eqn.2.20.
@@ 2.3.3 Error Analysis
From this point on, we drop the $t_{0}$ filtration for total expectation, variance and covariance operators by default unless explicitly specified, i.e, $\operatorname{\mathbb{E}}(\cdot)\equiv\operatorname{\mathbb{E}}(\cdot|\operatorname{\mathcal{F}}_{0})$, $\operatorname{\mathrm{Var}}(\cdot)\equiv\operatorname{\mathrm{Var}}(\cdot|\operatorname{\mathcal{F}}_{0})$ and $\operatorname{\mathrm{Cov}}(\cdot,\cdot\cdot)\equiv\operatorname{\mathrm{Cov}}(\cdot,\cdot\cdot|\operatorname{\mathcal{F}}_{0})$.
@@
$\sqrt{\frac{R-1}{R}(1-\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}_{k+1}^{1D}))}$, where $\tilde{\rho}_{k}(V_{k+1},\hat{V}_{k+1}^{1D})$ is defined in Eqn.2.34.
@@
We denote by $\Xi^{P^{\text{LSM}}}k[e^{-r\Delta T}V{k+1}]$ the upper bound of $\sqrt{\operatorname{\mathbb{E}}[(F^{P^{\text{LSM}}}_{k}-F_{k})^{2}]}$. A detailed derivation of this bound is provided in Appendix A.4. One of the key points is to express $F_{k}$ in terms of exact regression coefficients $\vec{\beta}^{P^{\text{LSM}}0}$ satisfying
@@
This result is consistent with previous studies on the asymptotic convergence rate of $\Xi^{P^{\text{LSM}}}_{k}$ going to zero under big $O_{p}$ notation by Stentoft, (2001) in their Theorem 1: the dominant term in the asymptotic from is related to the option variance and thus inversely proportional to $N_{R}$; whereas the other term $\sim R^{-\frac{2s}{d}}$ (with $s$ denoting the number of continuous derivatives of $F_{k}$ that exist) is negligible, i.e.,
@@
Apparently, the upper bound estimate using Eqn.2.37 is also valid for FD-LSM considering all the basis functions as a whole, but it is a loose one which ignores the correlation between $\vec{\phi}^{P^{\text{FD-LSM}}}_{R,k}$ and $V_{k+1}$. Thanks to the decomposition in Eqn.2.32 from Prop.1, we can achieve a tighter estimate as follows.
@@
$\begin{split}\operatorname{\mathbb{E}}[(F^{P^{\text{FD-LSM}}}_{k}-F_{k})^{2}]&=\operatorname{\mathbb{E}}[(P^{\text{FD}}_{k}[F^{P^{\text{FD-LSM}}}_{k}-F_{k}])^{2}]+\operatorname{\mathbb{E}}[(M^{\text{FD}}_{k}[F^{P^{\text{FD-LSM}}}_{k}-F_{k}])^{2}]\\ &=\operatorname{\mathbb{E}}[(M^{\text{FD}}_{k}[F^{P^{\text{FD-LSM}}}_{k}]-M^{\text{FD}}_{k}[F_{k}])^{2}].\end{split}$ (2.39)
Applying $M^{\text{FD}}_{k}$ to the right-hand side of Eqn.2.32 and exploiting the orthogonality condition $M^{\text{FD}}_{k}[F^{\text{FD}}_{k}[\cdot]]=0$, and then expressing $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$ via Eqn.2.25 and Eqn.2.27, we have
@@
$\begin{split}&\operatorname{\mathbb{E}}[(M^{\text{FD}}_{k}[F^{P^{\text{FD-LSM}}}_{k}]-M^{\text{FD}}_{k}[F_{k}])^{2}]\\ \leq&(R-1)\cdot\operatorname{\mathbb{E}}[(e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D})-(F_{k}-\beta_{k}^{\text{FD}}(F_{k})f^{\text{FD}}_{k}(X_{k})))^{2}]\\ \leq&(R-1)\cdot\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D})|\operatorname{\mathcal{F}}_{t_{k}})]\\ =&(R-1)[1-\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}^{1D}_{k+1})]\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(e^{-r\Delta T}V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}})],\end{split}$ (2.43)
where the identity $\operatorname{\mathbb{E}}[e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D})|\operatorname{\mathcal{F}}_{t_{k}}]=F_{k}-\beta_{k}^{\text{FD}}(F_{k})f^{\text{FD}}_{k}(X_{k})$ and Eqn.2.33 have been used in the penultimate and last line, respectively, apart from following the derivation described in Appendix A.4 with the mapping in Eqn.2.42.
@@
  • In the limit of $\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}^{1D}_{k+1})\to 0$, the control variate has null effect as $f^{\text{FD}}_{k}(X_{k})$ is orthogonal to $F_{k}$ in $L^{2}$. In this regime, the remaining mild error reduction, going as $\sqrt{\frac{R}{R-1}}$ is purely from reducing susceptibility of overfitting by removing unnecessary high order monomial component under the constraint of Eqn.2.38; further taking the limit of $R\to\infty$, $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$ converges to $P_{k}^{\text{LSM}}$ with MSE being equal to each other.
@@
The pricing error of FD-LSM converges to zero as $\tilde{\rho}_{m^{*}}(V_{m^{*}+1},\hat{V}^{1D}_{m^{*}+1})\to 1$, where $m^{*}:=\arg_{m\in[1,M]}\min\tilde{\rho}m(V{m+1},\hat{V}^{1D}_{m+1})$. This convergence holds in a generic multi-period (or multi-bucket) setting, while the error of the LSM counterpart remains finite. For the option PV problem, Assumption 1 is required; furthermore, if Assumption 2 is also satisfied, the convergence of the FD-LSM error remains valid even as $M\to\infty$.
@@
$\begin{split}\operatorname{\mathbb{E}}(V_{k}^{P}-V_{k})&=\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}((Z_{k}-V_{k+1})(\Theta(Z_{k}-P_{k}[V^{P}_{k+1}])-\Theta(Z_{k}-F_{k}))|\operatorname{\mathcal{F}}_{t_{k}}))+\\ &\qquad\operatorname{\mathbb{E}}((V_{k+1}^{P}-V_{k+1})\Theta(P_{k}[V^{P}_{k+1}]-Z_{k}))\\ &=\operatorname{\mathbb{E}}((Z_{k}-\operatorname{\mathbb{E}}(V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}}))(\Theta(Z_{k}-P_{k}[V^{P}_{k+1}])-\Theta(Z_{k}-F_{k})))+\\ &\qquad\operatorname{\mathbb{E}}((V_{k+1}^{P}-V_{k+1})\Theta(P_{k}[V^{P}_{k+1}]-Z_{k}))\\ &=\operatorname{\mathbb{E}}((Z_{k}-F_{k})(\Theta(Z_{k}-P_{k}[V^{P}_{k+1}])-\Theta(Z_{k}-F_{k})))+\\ &\qquad\operatorname{\mathbb{E}}(V_{k+1}^{P}-V_{k+1})\operatorname{\mathbb{E}}(\Theta(P_{k}[V^{P}_{k+1}]-Z_{k}))+\operatorname{\mathrm{Cov}}(V_{k+1}^{P}-V_{k+1},\Theta(P_{k}[V^{P}_{k+1}]-Z_{k})),\\ \end{split}$ (2.48)
@@
The problem is now reduced to analyzing the properties of an iterative backward sequence defined by $x_{k}=\sqrt{a_{k}(x_{k+1}^{2}+b)}+c_{k}x_{k+1}$, where $a_{k}=\mathcal{X}^{P}_{k},b=\Xi_{Max}^{P},c_{k}=\mathcal{Y}^{P}_{k}$, with the initial condition $x_{M}=0$.
@@
Next, we consider the asymptotic behavior of the error in the limit of $M\to\infty$ and $k\to 1$. As we can see, if Assumption 2 also holds, the fixed point condition of Eqn.4.10 in Appendix A.6.2 is satisfied with $a_{0}=\mathcal{X}^{P}_{0}$ and $c_{0}=\mathcal{Y}^{P}_{0}$. Consequently, using Eqn.4.9 and substituting $b=\Xi_{Max}^{P}$, one can see that the error saturates at
@@
as going backward. As usual, we have $\epsilon^{PV,P^{\text{FD-LSM}}}_{0}\sim O(\sqrt{1-\tilde{\rho}_{m^{*}}(V_{m^{*}+1},\hat{V}^{1D}_{m^{*}+1})})$ in the limit of $M\to\infty$.
@@
$\Xi_{Max}^{P^{\text{FD-LSM}}}=\left(5\%\times\sqrt{\frac{4-1}{4}(1-90\%^{2})}\right)^{2}=1.89\%^{2}$ as a numerical example along with $\mathcal{X}=0.01$ and $\mathcal{Y}=0.8$, we have the plots for $\epsilon^{PV,P}_{k}$ as depicted in Fig.2.1. The asymptotic limits obtained through backward iteration are consistent with the theoretical predictions of Eqn.2.54 2.54: specifically, $\epsilon^{PV,P^{\text{LSM}}}_{0}=2.89\%$ for LSM and $\epsilon^{PV,P^{\text{FD-LSM}}}_{0}=1.09\%$ for FD-LSM.
@@ 2.3.4 Implementation Details
Now we discuss the dimension reduction at multi-asset level. First of all, we can formally write down the effective 1D continuous dividend yield and local volatility as (see Avellaneda et al., (2002))

### v9

@@ 2.3.2 FD Solution: Control Variate
  • As mentioned before, $\hat{V}_{k+1}^{1D}$ can be seen as option value under a 1D simulation process. However, instead of employing the right-hand side in Eqn.2.33 directly for variance reduction, Prop.1 allows one to achieve exactly the same effect by just adding the FD ansatz into the projection basis functions. This helps mitigate computational cost as it saves additional tedious Monte Carlo simulation on $\hat{V}^{1D}$, as well as the variance and covariance calculation in Eqn.2.23 and 2.28.
  • From the proof of Prop.1, up to first order of variance reduction, we don’t see benefit of introducing cross term between $\vec{\phi}^{P^{\text{FD}}}(x)$ and $\vec{\phi}_{R-1}^{P^{\text{LSM}}}(x)$. This simplifies the construction of new basis and avoid overfitting due to further introduction of higher order terms.
  • The most effective FD auxiliary variable $\hat{V}^{1D}$ is obtained by achieving the highest possible correlation with $V$ based on Eqn.2.34. Heuristically, selecting a suitable ansatz can be seen as an effort to capture the dominant dynamics of the original high-dimensional problem within a computationally efficient and numerically tractable framework, e.g., 1D FD PDE, as an application of the Gyöngy theorem. This is done even if the maximal correlation of the resulting option price after dimension reduction is not explicitly proven. For instance, under a multi-asset single-factor setting (e.g. Black Scholes or Local Volatility), we can choose the ansatz from single-factor 1D PDE after inter-asset dimension reduction; for single-asset multi-factor settings (e.g. Heston stochastic volatility), one can construct the equivalent 1D process (Local volatility in general) by matching same marginal distributions as mentioned before in Sec.2.2.4; finally, for the most general case of multi-asset multi-factor setting, one could apply the dimension reduction on intra-asset level (reducing the factors other than spot, e.g. stochastic volatility on the same asset) to start with, and then on inter-asset level, eventually arriving at 1D problem which is solvable under Eqn.2.20.
@@ 2.3.3 Error Analysis
From this point on, we drop the $\operatorname{\mathcal{F}}_{0}$ filtration for total expectation, variance and covariance operators by default unless explicitly specified, i.e, $\operatorname{\mathbb{E}}(\cdot)\equiv\operatorname{\mathbb{E}}(\cdot|\operatorname{\mathcal{F}}_{0})$, $\operatorname{\mathrm{Var}}(\cdot)\equiv\operatorname{\mathrm{Var}}(\cdot|\operatorname{\mathcal{F}}_{0})$ and $\operatorname{\mathrm{Cov}}(\cdot,\cdot\cdot)\equiv\operatorname{\mathrm{Cov}}(\cdot,\cdot\cdot|\operatorname{\mathcal{F}}_{0})$.
@@
$\sqrt{\frac{R-1}{R}(1-\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}_{k+1}^{1D}))}$, where $\tilde{\rho}_{k}(V_{k+1},\hat{V}_{k+1}^{1D})$ is defined in Eqn.2.35.
@@
We denote by $\Xi^{P^{\text{LSM}}}_{k}[e^{-r\Delta T}V_{k+1}]$ the upper bound of $\sqrt{\operatorname{\mathbb{E}}[(F^{P^{\text{LSM}}}_{k}-F_{k})^{2}]}$. A detailed derivation of this bound is provided in Appendix A.4. One of the key points is to express $F_{k}$ in terms of exact regression coefficients $\vec{\beta}^{P^{\text{LSM}}0}$ satisfying
@@
This result is consistent with previous studies on the asymptotic convergence rate of $\Xi^{P^{\text{LSM}}}_{k}$ going to zero under big $O_{p}$ notation by Stentoft, (2001) in their Theorem 1: the dominant term $R/N_{R}$ in the asymptotic from is related to the option variance and thus inversely proportional to $N_{R}$; whereas the other term $R^{-\frac{2s}{d}}$ (with $s$ denoting the number of continuous derivatives of $F_{k}$ that exist) is negligible, i.e.,
@@
Apparently, the upper bound estimate using Eqn.2.37 is also valid for FD-LSM considering all the basis functions as a whole, but it is a loose one which ignores the correlation between $\vec{\phi}^{P^{\text{FD-LSM}}}_{R,k}$ and $V_{k+1}$. Thanks to the decomposition in Eqn.2.33 from Prop.1, we can achieve a tighter estimate as follows.
@@
$\begin{split}\operatorname{\mathbb{E}}[(F^{P^{\text{FD-LSM}}}_{k}-F_{k})^{2}]&=\operatorname{\mathbb{E}}[(P_{k}^{\text{FD}}[F^{P^{\text{FD-LSM}}}_{k}-F_{k}])^{2}]+\operatorname{\mathbb{E}}[(M^{\text{FD}}_{k}[F^{P^{\text{FD-LSM}}}_{k}-F_{k}])^{2}]\\ &=\operatorname{\mathbb{E}}[(M^{\text{FD}}_{k}[F^{P^{\text{FD-LSM}}}_{k}]-M^{\text{FD}}_{k}[F_{k}])^{2}].\end{split}$ (2.39)
Applying $M^{\text{FD}}_{k}$ to the right-hand side of Eqn.2.33 and exploiting the orthogonality condition $M^{\text{FD}}_{k}[f^{\text{FD}}_{k}(X_{k})]=f^{\text{FD}}_{k}(X_{k})-P^{\text{FD}}_{k}[f^{\text{FD}}_{k}(X_{k})]=0$, and then expressing $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$ via Eqn.2.25 and Eqn.2.27, we have
@@
$\begin{split}&\operatorname{\mathbb{E}}[(M^{\text{FD}}_{k}[F^{P^{\text{FD-LSM}}}_{k}]-M^{\text{FD}}_{k}[F_{k}])^{2}]\\ \leq&(R-1)\cdot\operatorname{\mathbb{E}}[(e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D})-(F_{k}-\beta_{k}^{\text{FD}}(F_{k})f^{\text{FD}}_{k}(X_{k})))^{2}]\\ =&(R-1)\cdot\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D})|\operatorname{\mathcal{F}}_{t_{k}})]\\ =&(R-1)[1-\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}^{1D}_{k+1})]\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(e^{-r\Delta T}V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}})],\end{split}$ (2.43)
where the identity $\operatorname{\mathbb{E}}[e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D})|\operatorname{\mathcal{F}}_{t_{k}}]=F_{k}-\beta_{k}^{\text{FD}}(F_{k})f^{\text{FD}}_{k}(X_{k})$ and Eqn.2.34 have been used in the penultimate and last line, respectively, apart from following the derivation described in Appendix A.4 with the mapping in Eqn.2.42.
@@
  • In the limit of $\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}^{1D}_{k+1})\to 0$, the control variate has null effect as $f^{\text{FD}}_{k}(X_{k})$ is orthogonal to $F_{k}$ in $L^{2}$. In this regime, the remaining mild error reduction, going as $\sqrt{\frac{R-1}{R}}$ is purely from reducing susceptibility of overfitting by removing unnecessary high order monomial component under the constraint of Eqn.2.38; further taking the limit of $R\to\infty$, $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$ converges to $P_{k}^{\text{LSM}}$ with MSE being equal to each other.
@@
The pricing error of FD-LSM converges to zero as $\tilde{\rho}_{m^{*}}(V_{m^{*}+1},\hat{V}^{1D}_{m^{*}+1})\to 1$, where $m^{*}:=\arg_{m\in[1,M]}\min\tilde{\rho}_{m}(V_{m+1},\hat{V}^{1D}_{m+1})$ represents the time step at which the minimum correlation between $V$ and $\hat{V}^{1D}$ is achieved. This convergence holds in a generic multi-period (or multi-bucket) setting, while the error of the LSM counterpart remains finite. For the option PV problem, Assumption 1 is required; furthermore, if Assumption 2 is also satisfied, the convergence of the FD-LSM error remains valid even as $M\to\infty$.
@@
$\begin{split}\operatorname{\mathbb{E}}(V_{k}^{P}-V_{k})&=\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}([Z_{k}-V_{k+1}][\Theta(Z_{k}-P_{k}[V^{P}_{k+1}])-\Theta(Z_{k}-F_{k})]|\operatorname{\mathcal{F}}_{t_{k}}))+\\ &\qquad\operatorname{\mathbb{E}}[(V_{k+1}^{P}-V_{k+1})\Theta(P_{k}[V^{P}_{k+1}]-Z_{k})]\\ &=\operatorname{\mathbb{E}}([Z_{k}-\operatorname{\mathbb{E}}(V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}})][\Theta(Z_{k}-P_{k}[V^{P}_{k+1}])-\Theta(Z_{k}-F_{k})])+\\ &\qquad\operatorname{\mathbb{E}}[(V_{k+1}^{P}-V_{k+1})\Theta(P_{k}[V^{P}_{k+1}]-Z_{k})]\\ &=\operatorname{\mathbb{E}}((Z_{k}-F_{k})[\Theta(Z_{k}-P_{k}[V^{P}_{k+1}])-\Theta(Z_{k}-F_{k})])+\\ &\qquad\operatorname{\mathbb{E}}(V_{k+1}^{P}-V_{k+1})\operatorname{\mathbb{E}}(\Theta(P_{k}[V^{P}_{k+1}]-Z_{k}))+\operatorname{\mathrm{Cov}}[(V_{k+1}^{P}-V_{k+1}),\Theta(P_{k}[V^{P}_{k+1}]-Z_{k})],\\ \end{split}$ (2.48)
@@
The problem now lies in analyzing the characteristics of an iterative backward sequence, governed by the equation $x_{k}=\sqrt{a_{k}(x_{k+1}^{2}+b)}+c_{k}x_{k+1}$. The (non-negative) parameters are defined as follows: $a_{k}=\mathcal{X}^{P}_{k}$, $b=\Xi_{Max}^{P}$, and $c_{k}=\mathcal{Y}^{P}_{k}$. This sequence is initialized with the condition $x_{M}=0$. A detailed analysis of this sequence’s properties is presented in Appendix A.6.
@@
Next, we consider the asymptotic behavior of the error in the limit of $M\to\infty$ and $k\to 1$. As we can see from the fixed point analysis (Appendix A.6.2) of the backward sequence, if Assumption 2 also holds, the fixed point condition of Eqn.4.10 is satisfied with $a_{0}=\mathcal{X}^{P}_{0}$ and $c_{0}=\mathcal{Y}^{P}_{0}$. Consequently, using Eqn.4.9 and substituting $b=\Xi_{Max}^{P}$, one can see that the error saturates at
@@
as going backward. From above, we arrive at $\epsilon^{PV,P^{\text{FD-LSM}}}_{0}\sim O(\sqrt{1-\tilde{\rho}_{m^{*}}(V_{m^{*}+1},\hat{V}^{1D}_{m^{*}+1})})$, similar to the case in Eqn.2.53 with finite $M$.
@@
$\Xi_{Max}^{P^{\text{FD-LSM}}}=\left(5\%\times\sqrt{\frac{4-1}{4}(1-90\%^{2})}\right)^{2}=1.89\%^{2}$ as a numerical example along with $\mathcal{X}=0.01$ and $\mathcal{Y}=0.8$, we have the plots for $\epsilon^{PV,P}_{k}$ as depicted in Fig.2.1. The asymptotic limits obtained through backward iteration are consistent with the theoretical predictions of Eqn.2.54: specifically, $\epsilon^{PV,P^{\text{LSM}}}_{0}=2.89\%$ for LSM and $\epsilon^{PV,P^{\text{FD-LSM}}}_{0}=1.09\%$ for FD-LSM.
@@ 2.3.4 Implementation Details
Now we discuss the dimension reduction at multi-asset level. First of all, we can formally write down the effective 1D continuous dividend yield and local volatility as (see Avellaneda et al., (2002); Gyöngy, (1986))